# Feedback Connections in Quantum Reservoir Computing with Mid-Circuit Measurements

### v1

@@ IV Results and discussion
We obtain all predictive and memory related results by simulating a minimal instance of the proposed architecture with a system size of $N=2$. While we do use a density matrix formalism, accounting for all possible measurement trajectories under the conditional feedforward mechanism would require exponential overhead and is therefore not feasible. Instead, we employ a shot-based approach, in which each shot represents a single experiment on quantum hardware, and individual measurement outcomes are sampled directly. For further computational efficiency, we choose a washout length $l_{w}=25$, training length $l_{\mathrm{tr}}=100$, and test length $l_{\mathrm{ts}}=100$.
@@
Figure 3: Evaluation of short term memory and predictive capabilities: All values are averaged over 128 Haar-random unitaries. Since there is no single ‘privileged’ choice of unitary, our primary interest is the mean performance of the protocol over these random instances. We therefore report the standard error of the mean to reflect how much this mean would vary if we resampled from the Haar measure many times.(a) The left panel shows the short-term memory (STM) capacity for several different values of $a_{fb}$, with $\tau$ ranging from 00 to $-3$. The right panel provides a more detailed view for $a_{fb}=1.3$, plotting all values of $\tau$. (b) Here we display the NMSE for the quantum 1D Ising chain, showing predictive capabilities for $\tau=1$ across various $a_{fb}$ values. The right panel presents predictive capabilities for different $\tau$ values for selected $a_{fb}$ values. Figure (c) shows the same analysis for the Mackey-Glass task. For both (b) and (c), $5k$ shots were used.
@@
In Fig. 3 (a) on the right, we present the short-term memory correlation coefficient for four different delay values. Even with a limited number of shots, the minimal example exhibits nonzero correlation coefficients for $\tau=1$, indicating that the fundamental feedback connections during continuous processing play a beneficial role for memory. Notably, the feedback connections were evaluated using a reset operation after each measurement, ensuring that any information transmitted forward originates solely from the proposed mechanism. Furthermore, an upward shift is observed for an increasing number of shots, suggesting that performance improves with additional measurement statistics. On the left in 3 (a), we show the effect of varying the feedback strength on the memory capacity $C_{\Sigma}$ for different numbers of measurements for a fixed $a_{fb}=1.3$. We observe again an upward shift in performance with an increasing number of measurement shots. Notably we think that careful tuning of both the feedback and input strengths is crucial for robust performance as well as the right choice of the Haar matrix. As noted in [14], when the angles in the feedback and input $R$ gates exceed ±$\pm 2\pi$, the reservoir fails to process the feedback connections reliably, causing the memory capacity to vanish. From the evaluation in Fig. 3 (a) we can see, that there is an increase in memory capacity for $a_{fb}$ values from $1.0$ to around $2.0$. Interestingly, this trend is also reflected in the predictive performance — lower NMSE values are observed for the same $a_{fb}$ range, as illustrated in Fig. 3 (b) and (c). This again affirms the importance of memory capacity for predictive tasks. We can report an overall minimal average NMSE of $8.8\times 10^{-3}$ on the 5-qubit Ising chain. The model performs slightly worse on the classical dynamical MG sequence with a minimal average NMSE of $1.9\times 10^{-2}$.
@@
Figure 4: Echo state property for various architectures: The top figures show the first element of the reservoir state over time for five randomly initialized runs, while the bottom figures present the absolute difference among these five reservoir states. (a) Classical ESN model. (b) Feedback driven QRC regime [14]. (c) Mid-circuit measurement QRC [20]. (d) Our algorithm.
@@
We now present the results from the experiment on the analysis of the echo state property of different reservoir computing models as described in subsection III-C. In all methods, we replicate the reservoirs as described in their respective references. In particular:
  • For the classical minimal ESN model - we use a network of the form
@@
    where $x_{k}$ contains 1000 dimensions, each element in $W_{in}$ , $b$ and $W$ are initialized with the random variable $X\sim\mathcal{N}(-0.5,1)$ in addition to which $W$ is then re-normalized to have a spectral radius of 1.25, and $\alpha=0.3$. This configuration replicates a model that was able to accurately reproduce the Mackey-Glass trajectory with a delay value $\tau$ of 17 in an auto-regressive manner for over 2000 steps after the training cycle. Each run was performed with a different initializaion of $x_{0}$, with each parameter randomly chosen from a uniform distribution between 0 and 1.
  • For the feedback driven QRC protocol - we take the architecture proposed in Fig. 1 (c) in Ref. [14] with the hyperparameters $a_{in}=0.001$, $a_{fb}=2.5$ and a Haar-randomly chosen $U_{res}$, and using a statevector simulator to compute the expectation values of the measurements. This configuration replicates the model used to produce the results in the reference. Each run was performed with a different initialization of the initial feedback vector, with each component randomly chosen from a uniform distribution between 0 and 1.
  • For the QRC scheme with a mid-circuit measurement architecture - we take the architecture proposed in Fig. 1 (d) in Ref. [20] with 10 time-steps and setting $a=5$ and $10^{4}$ shots222although no specific value for $a$ was explicitly stated in the original work, we were able to reproduce the NARMA2 benchmark results when setting $a=5$. This model requires no random initializations.
  • For the model proposed in this paper - we use the model described in Fig. 1 (e) with $a_{in}=1$, $a_{fb}=1.6$, a Haar-randomly chosen $U_{\text{Haar}}$ and $10^{4}$ shots. This configuration maximized the short-term memory capacity as shown in Fig 3 (a). Each run was performed with a different initialization of the initial random feedback vector $\textbf{m}_{-1}$.
The results of the ESP analysis can be found in Fig. 4. We notice that the classical minimal ESN model and the feedback driven QRC protocol model both clearly display the echo-state property, see Fig. 4 (a) and (b). On the other hand, we see that the QRC scheme with a mid-circuit measurement architecture and the model we propose do not display reservoirs that are asymptotically independent of the initial state. However, the QRC scheme with a mid-circuit measurement architecture, despite not having an echo state property, was shown to be successful at reproducing NARMA time-series benchmarks and to maybe even have a promise in modeling soft robots [20]. Therefore the relationship between the asymptotic behavior of a quantum reservoir state with respect to its initial conditions vs the reservoir model having the fading-memory property does not seem to be as straightforward for the QRC models as it is for their classical counterparts. To ensure we did not happen to fall upon a special case, we repeated the analysis with several different initializations of the fixed random parameters (e.g. $W_{in}$, $U_{\text{Haar}}$, etc.) and across different dimensions of the internal state of the reservoir, which led to qualitatively similar results as the ones shown in Fig. 4.
@@ V Conclusion
In this paper, we propose a quantum reservoir computing scheme that uses mid-circuit measurements and feedback within the runtime time of a qubit. Our initial experiments aim to determine whether feedback connections during runtime genuinely improve memory capacity rather than merely introducing uncertainties due to measurement noise. We show that an architecture of this type with incorporated resets exhibits nonzero memory capacity for past inputs, indicating that feedback connections have an impact. Notably, peaks in memory capacity at specific feedback strengths align with enhanced performance on predictive tasks involving both classical and quantum chaotic systems. Our ESP findings reveal that while the classical minimal ESN model and the feedback-driven QRC protocol both exhibit clear echo-state behavior, mid-circuit measurements QRC schemes can remain effective despite lacking this property. Nonetheless, we strongly think that this idea is yet not fully explored. Further investigation could involve experimenting with alternative reset strategies, such as using conditional resets or omitting resets entirely, thus allowing the model to continue its evolution from the post-measurement state, while additionally having feedback connections. We have also not yet explored scaling the system size, which is an is promising direction for future investigation that can unveil scaling capacities of this approach. Finally, the most important next step is to implement this approach on actual quantum hardware. As noted in Section I, the IBM platform supports mid-circuit measurements with classical feedforward operations, but these are currently limited to 50 operations per circuit. This constraint will likely be eased in the future, particularly as error-correcting protocols become more prevalent. We would also like to suggest to dive deeper in the underlying importance of the Haar random unitary in QRC schemes.
@@ Appendix A Noise Test
Here we present some preliminary noise tests. In our numerical shot based simulation we introduce a depolarization channel after each unitary of the form:
@@
where $\lambda$ is the probability of the state to be found in the completely mixed state $\frac{I}{d}$. For the one qubit case, the DP channel can expressed also with the Kraus operators
@@
We set lambda to $0.04$, which corresponds geometrically to a uniformly Bloch sphere reduction of $4\%$, or in terms of random processes: With $97\%$ chance, we apply the identity. With $1\%$ chance each, we apply $X$, $Y$ or $Z$. We now look at the predictive capabilities on the Ising chain, similar to Fig. 3 (b).
This preliminary finding suggests that our architecture exhibits noise resistance. Investigation will go into the explanation of why this is the case.
@@ Appendix B Code Availability
The Python implementation to reproduce the results can be found at https://github.com/Muri7910/Feedback-Connections-in-Quantum-Reservoir-Computing-with-Mid-Circuit-Measurements. Further information and data is available upon reasonable request.

### v3

@@ IV Results and discussion
In this section, we report on three key topics: simulation results of our proposed model, a hardware implementation on a QPU, and a simulation-based echo state property analysis for several reservoir computing models.
IV-A Memory capacity and predictive performance
@@
We obtain all predictive and memory results by simulating a minimal instance of the proposed architecture with a system size of $N=2$. We use a shot-based approach, where each shot represents a single experiment on the quantum hardware, and individual measurement results are sampled directly. This approach mimics a noiseless version of what is done on real hardware, and allows us to take into account the effect of shot noise which plays an important role in the performance of quantum reservoir models [32]. We also note that since we use a density matrix formalism, analytically considering all possible measurement trajectories under the conditional feedforward mechanism would require exponential overhead and is therefore not feasible. For further computational efficiency, we choose a washout length $l_{w}=25$, a training length $l_{\mathrm{tr}}=100$, and a test length $l_{\mathrm{ts}}=100$.
@@
Figure 3: Evaluation of short term memory and predictive capabilities: All values are averaged over 128 Haar-random unitaries. Since there is no single ‘privileged’ choice of unitary, our primary interest is in the average performance of the protocol over these random instances. Therefore we choose to indicate one standard deviation of the mean with the error bars to show how much this mean would vary if we resampled the Haar measure many times. (a) The left panel shows the short-term memory (STM) capacity for several different values of $a_{fb}$, with $\tau$ ranging from 00 to $-3$. The right panel provides a more detailed view for $a_{fb}=1.3$, plotting all values of $\tau$. (b) Here, we plot the NMSE for the quantum 1D Ising chain, showing the prediction capabilities for $\tau=1$ over different $a_{fb}$ values. The right panel shows the prediction capabilities for different $\tau$ values for selected $a_{fb}$ values. (c) shows the same analysis for the Mackey-Glass task. $5k$ shots were used for the right panels in both (b) and (c).
@@
In Fig. 3 (a) on the right, we plot the short-term memory correlation coefficient for four different delay values for a fixed $a_{fb}=1.3$. Even with a limited number of shots, the minimal example shows nonzero correlation coefficients for $\tau=-1$, indicating that the basic feedback connections during continuous processing play a beneficial role for memory. Notably, the feedback connections were evaluated using a reset operation after each measurement, ensuring that any information carried forward is solely from the proposed mechanism. Furthermore, an upward shift is observed for an increasing number of shots, suggesting that performance improves with additional measurement statistics. On the left in Fig. 3 (a) we show the effect of varying the feedback strength on the memory capacity $C_{\Sigma}$ for different numbers of measurements. Again, we observe an upward shift in performance with increasing number of measurements. In particular, we believe that careful tuning of both feedback and input strengths is crucial for robust performance, as is the right choice of Haar matrix. As noted in [9], if the angles in the feedback and input $R$ gates exceed ±$\pm 2\pi$, the reservoir cannot reliably process the feedback connections and the memory capacity disappears. From the analysis in Fig. 3 (a) we can see that there is an increase in memory capacity for $a_{fb}$ values from $1.0$ to about $2.0$. Interestingly, this trend is also reflected in the prediction performance – lower NMSE values are observed for the same $a_{fb}$ range, as shown in Fig. 3 (b) and (c). This again confirms the importance of memory capacity for predictive tasks. We can report an overall minimum average NMSE of $5.07\times 10^{-3}$ on the 5-qubit Ising chain. The model performs slightly worse on the classical dynamical Mackey-Glass sequence with a minimum average NMSE of $1.4\times 10^{-2}$. We also note that the model is robust to some amount of depolarizing noise (more details in Appendix A).
@@
Figure 4: Echo state property for different architectures: The top figures show the first component of the reservoir state over time for five randomly initialized runs, while the bottom figures show the absolute difference between these five reservoir states. (a) Classical ESN model. (b) Feedback-driven QRC regime [9]. (c) Mid-circuit measurements QRC [15]. (d) Proposed feedback-based QRC with mid-circuit measurements and post-measurement resets.(e) Proposed feedback-based QRC with mid-circuit measurements and continuous operation (no resets).
IV-B QPU results
We now present experiments with our architecture on IBM hardware, specifically utilizing the dynamic circuit capabilities of ibm_marrakesh. Due to hardware constraints on the number of resets and conditional operations, we reduced the washout, training, and testing lengths to 5, 25, and 20 respectively. Additionally, time and hardware limitations made it infeasible to average over 128 Haar-random unitaries, so we selected a single fixed unitary. Below, we briefly describe how we generate a Haar-random unitary for the IBM device. A more detailed description of this algorithm is found in [33] and [34]:
  1. 1.
    Generate a $2^{N}\times 2^{N}$ matrix $A$ with complex entries $z_{i}=a_{i}+b_{i}i$, where all $a_{i},b_{i}$ are sampled from the Ginibre ensemble
  2. 2.
    Compute the QR decomposition $A=QR$
  3. 3.
    Compute the diagonal $\Lambda=\text{diag}(R_{ii}/|R_{ii}|)$
  4. 4.
    Compute $A^{\prime}=Q\Lambda$, where $A^{\prime}$ is a Haar-random unitary
We then let the qiskit [35] compiler find native gates that best approximates this unitary $A^{\prime}$.
Figure 5: QPU Test: NMSE for various $\tau$ using (a) Ising chain and (b) Mackey-Glass sequences.
In Fig. 5, we compare the NMSE of our QPU experiment with that of the simulation, both run with $10^{4}$ shots on the Ising chain and Mackey-Glass sequence. We note that both the simulation and the QPU runs have inherently limited predictive capabilities (as seen in the very high NMSE values) because of the greatly reduced training and test lengths. Nonetheless, our main objective here is to demonstrate that the proposed method can be executed on current quantum hardware.
IV-C Echo state property analysis
@@
We now present the results of the experiment on the analysis of the echo state property of different reservoir computing models. In all methods, we replicate the reservoirs as described in the respective references. In particular:
  • For the classical minimal ESN model [31] – we use a network of the form:
@@
    where $x_{k}$ is 1000-dimensional. Each element in $W_{\mathrm{in}},b,$ and $W$ is initialized with the random variable $X\sim\mathcal{N}(-0.5,1)$. Next, $W$ is renormalized so that its spectral radius is 1.25, and $\alpha$ is set to 0.3. This configuration reproduces a model that performed well on the Mackey–Glass sequence used throughout this paper. Each run was performed with a different initialization of $x_{0}$, with each component randomly drawn from a uniform distribution between 0 and 1.
  • For the feedback-driven QRC protocol [9] – we take the architecture proposed in Fig. 1 (c) with the hyperparameters $a_{in}=0.001$, $a_{fb}=2.5$ and a Haar-randomly chosen $U_{res}$, using a state vector simulator to compute the expectation values of the measurements. Each run was performed with a different Haar-random initialization of the start state.
  • For the QRC scheme with a mid-circuit measurement architecture [15] – we take the architecture proposed in Fig. 1 (d) with 10 time steps and setting $a=5$ and $10^{4}$ shots [^2] Although no specific value for $a$ was explicitly given in the original work, we were able to reproduce similar results to original work’s NARMA2 benchmark results when setting $a=5$.. Each run was performed with a different Haar-random 2-qubit initial state while keeping the 2 ancilla qubits always starting in the |$\ket{0}$ state.
  • For the model proposed in this paper – we use the model described in Fig. 1 (e) with $a_{in}=1$, $a_{fb}=1.6$, a Haar-randomly chosen $U_{\text{Haar}}$ and $10^{4}$ shots. Each run was performed with a different Haar-random 2-qubit initial state (denoted by $\rho_{0}$ in Fig. 1 (e)).
The results of the ESP analysis are shown in Fig. 4. We see that all models aside from the proposed model with post-measurement resets, clearly exhibit the echo-state property. And even for the proposed model with post-measurement resets, we see that effect of the initial state becomes less prominent over time, although it does not completely vanish. This shows that all the QRC models investigated exhibit ESPs to at least some extent. To make sure that we did not stumble upon a special case, we found qualitatively similar results when we repeating the analysis with several different initializations of the fixed random parameters (e.g., $W_{in}$, $U_{\text{Haar}}$, etc.) and over different dimensions of the internal state of the reservoir.
@@ V Conclusion
In this paper, we propose a quantum reservoir computing scheme that uses mid-circuit measurements and feedback within the runtime of a qubit. Our initial experiments aim to determine whether feedback connection that happen during the coherence time of the qubit actually improves memory capacity, rather than merely introducing uncertainties due to measurement noise. We show that such an architecture with built-in resets has non-zero memory capacity for past inputs, indicating that feedback connections have an impact. In particular, peaks in memory capacity at specific feedback strengths are consistent with improved performance on prediction tasks involving both classical and quantum chaotic systems. Our ESP results show that most investigated QRC regimes exhibit the ESP, while our proposed model with post-measurement resets seems to have a mild form of the ESP. Further investigations are warranted to understand the importance of the ESP on the performance of QRC models.
This work demonstrates sufficiently encouraging results to warrant further investigations into the approach of having feedback connections during coherence time. Given that our model exhibits some initial memory and prediction capabilities even at very small scales, exploring the scaling of our model is a very promising direction for future investigation that can reveal more on the effectiveness of this approach. Further investigations could also include experimenting with alternative reset strategies, such as using conditional resets or omitting resets altogether, thus allowing the model to continue to evolve from the post-measurement state while also having feedback connections.
@@ Appendix A Noise Test
We present preliminary noise tests for larger runs, accounting for the significant reductions in QPU runs. We use the same run statistics as in our numerical experiments. In our shot-based simulation, we introduce a depolarization channel after each unitary of the form:
@@
where $\lambda$ is the probability of the state to be found in the completely mixed state $\frac{I}{d}$. For the one qubit case, the DP channel can expressed also with the Kraus operators:
@@
We set lambda to $0.04$, which corresponds geometrically to a uniform Bloch sphere reduction of $4\%$, or in terms of random processes: With $97\%$ chance we apply the identity. With $1\%$ chance we use $X$, $Y$ or $Z$. The predictive capabilities shown in Fig. 6 on the Ising chain (similar to Fig. 3 (b)) demonstrate that our model is robust to some amount of depolarizing noise.
Figure 6: Noise Test: We show the noise effect of on our QRC regime on the NMSE of the Ising chain prediction tasks.
@@ Appendix B Code Availability
The Python implementation can be found at https://github.com/Muri7910/FB_Con_MCM_QRC.